Q . NO . 1 ( 25 ) ½ x ( 32 ) 2/5 divide ( 27 ) 2/3 x ( 16 ) 1/4

Q .  NO . 1    
( 25 ) ½    x   ( 32 ) 2/5    
                           divide  
                   ( 27 ) 2/3     x    (  16 ) 1/4
 

Grade:9

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given
=\frac{25^{1/2}32^{2/5}}{27^{2/3}16^{1/4}}
=((52)1/2(25)2/5)/(33)2/3(24)1/4
=(5*4)/(9*2)
=20/18
=10/9